Show that the set W = {(a, b, 0) : a, b ∈ F} is a subspace of V3(F).

1
Expert's answer
2022-01-17T07:00:01-0500

Let us show that the set "W = \{(a, b, 0) : a, b ∈ F\}" is a subspace of "V_3(F)."

Let "(a, b, 0),(c, d, 0)\in W."

Then "(a, b, 0)+(c, d, 0)=(a+c,b+d,0)\in W."

If "f\in F," then "f\cdot(a,b,0)=(fa,fb,0)\in W."

We conclude that "W = \{(a, b, 0) : a, b ∈ F\}" is a subspace of "V_3(F)."